The lifetime of a microprocessor is exponentially distributed with a variance of 4,000,000 hours.

a. What proportion of microprocessors will function for less than 5,000 hours?

b. A microprocessor has been functioning for 1,000 hours. What is the probability that it will

function for a total of at least 6,000 hours?
Answer #1

Ans:

a)

mean=sqrt(4000000)=2000

P(T<=5000)=1-exp(-5000/2000)=0.9179

b)

P(T>=6000/T>=1000)=P(T>=5000)

=exp(-5000/2000)=0.0821